OpenStudy (anonymous):

how much will $850 amount to be in three years if it is invested at 8% interest compounded quarterly for 3 years? A. $886.47 B. $952.43 C. $1078.01 D. $1168.15

OpenStudy (mathstudent55):

Compound Interest Formula \(\large F = P(1 + \dfrac{i}{n})^{nt}\) F = future value P = present value i = annual interest rate written as a decimal n = number of times compounded per year t = number of years

OpenStudy (mathstudent55):

No. Be careful how you plug in the numbers in the formula. i = 8%, so use i = 0.08 Also, compounded quarterly means 4 times per year, so n = 4 t is the number of years, so t = 3

OpenStudy (fateal):

so F= 850(1+0.08/4)^8*4? can you just help me set it up the right way? thank you

OpenStudy (mathstudent55):

\(\large F = $850 \left(1 + \dfrac{0.08}{4} \right)^{4 \times 3} \)

OpenStudy (mathstudent55):

You almost have it now. Notice your exponent. You have 8 * 4 8 is the interest rate in percent. It does not go on the exponent. The exponent is the number of compoundings per year (4) times the number of years(3)

OpenStudy (fateal):

so the answer is C? thank you ^^

OpenStudy (mathstudent55):

Yes, answer is C.